# Hierarchical Reinforcement Learning for Swarm Confrontation with High Uncertainty

### v1

@@ Abstract
In swarm robotics, confrontation including the pursuit–evasion game is a key scenario. High uncertainty caused by unknown opponents’ strategies and dynamic obstacles complicates the action space into a hybrid decision process. Although the deep reinforcement learning method is significant for swarm confrontation since it can handle various sizes, as an end–to–end implementation, it cannot deal with the hybrid process. Here, we propose a novel hierarchical reinforcement learning approach consisting of a target allocation layer, a path planning layer, and the underlying dynamic interaction mechanism between the two layers, which indicates the quantified uncertainty. It decouples the hybrid process into discrete allocation and continuous planning layers, with a probabilistic ensemble model to quantify the uncertainty and regulate the interaction frequency adaptively. Furthermore, to overcome the unstable training process introduced by the two layers, we design an integration training method including pre–training and cross–training, which enhances the training efficiency and stability. Experiment results in both comparison and ablation studies validate the effectiveness and generalization performance of our proposed approach.
@@ I Introduction
Swarm confrontation, being a typical complex problem solving, naturally consists of discrete and continuous spaces under an uncertain environment. Illustratively, commands on the battlefield always exist in the form of discrete decisions, while actions usually take place in continuous time. Recent works [24, 25] combine commands and actions into multiple high–level spaces. This method blurs the interpretability of the spaces inevitably resulting in slowly converging algorithms for large–scale swarms. Alternatively, it is not hard to reflect the swarm confrontation to divide–and–conquer framework, where the commands are translated into target allocation and the actions are addressed as path planning, from the artificial intelligence perspective[10]. Wang et al. [26] make an interesting attempt for the first time to introduce HRL into swarm confrontation. Inspired by the method, [27] decomposes multi–aircraft formation air combat into high–level strategy and sub–strategy, achieving favorable effects in confrontation among a few robots. However, the fact that we cannot guarantee the stability of the algorithms prevents us from building a bridge between HRL and swarm confrontation. Others [28, 19] have already improved the stability of the algorithm for HRL in a different scenario. They design an interactive training strategy, including pre–training, intensive training, and alternate training, to ensure the stability guaranteed. A prerequisite for this strategy is only the lower layer pre–trained, while the training of the upper is limited by the cumulative feedback from the lower. Thus we cannot directly implement it to swarm confrontation for the environment uncertainty will make the global optimization impossible.
@@
Here, we propose a guaranteed stable HRL method, which induces quantified uncertainty into an interaction mechanism linking the allocation and planning layers, to solve the hybrid problem of swarm confrontation in various sizes and environments. Firstly, we construct two–layer DRL networks to reflect commands and actions into target allocation and path planning, respectively, since the high uncertainty caused by the nature of the confrontation, including variant opponents’ strategies and transient battlefield environment, demands a hybrid, flexible, and robust intelligent algorithm. Secondly, the mechanism, which is embedded with a probabilistic ensemble model [29] quantifying the uncertainty, regulates the interaction frequency between the two layers. The essence is the frequency is increased as the circumstance becomes uncertain and dire. Thirdly, a novel integration training method (ITM), consisting of pre–training and cross–training, ensures the stability of HRL, in which, notably, we combine pre–trained and an improved model–based value estimation (IMVE) [30] method together to fasten the convergent speed of the upper in case of few samples given by the lower. Finally, extensive experiments on different–size swarms verify that our approach outperforms the baselines including non–learning approach and traditional DRL, and they also demonstrate the necessity of adaptive frequency approach and ITM through ablation studies. Plus, our method shows that a trained model under a small scale holds favorable generalization in various scales of swarm confrontation. The main contributions of this paper are summarized as follows.
  • We propose a novel HRL framework that enables multiple agents to intelligently interact with the environment under high uncertainty, overcoming the obstacle that traditional DRL cannot handle disturbing hybrid spaces fluently.
  • The adaptive interaction mechanism proposed in HRL eliminates the high uncertainty caused by some special scenarios like intelligent confrontation and the integration training method guarantees the whole system stable.
  • Our method solves the realization problem of swarm confrontation, demonstrating a new way for artificial intelligence to be applied in both civil and military scenes.
@@
We organize the rest of the paper as follows. Section II introduces the research related to our study. Section III presents the problem formulation of our problem. Section IV presents our two–layer networks and guaranteed stable HRL method. Section V describes the experiments of our method. Section VI presents our conclusions.
@@ II Related Works
The expert system method [10] approaches the system with a prior knowledge of human experts and selects the strategies in the knowledge base by fuzzy matching method. Wang et al.[31] develop an effective self–organized swarm confrontation decision–making method consisting of task allocation and swarm motion. The method relies on rules developed by an enormous number of human experts and is unable to ensure the optimality of decisions in a complex swarm confrontation environment. In game theory[11], the swarm confrontation problem is frequently modeled as a differential game. Liu et al.[32] adopt the Nash equilibrium seeking strategy to respond to agents in a dual–coalition non–cooperative game. Game theory suffers from problems such as too many state variables and complex differential equations, making it difficult to apply to complex multi–agent environments. The heuristic approach [3] considers modeling the swarm as biologically inspired networks to simulate the dynamics of swarm confrontation. Liu et al.[33] propose a novel evolutionary algorithm–based attack strategy with swarm robots, which has more potential to solve large–scale confrontational problems. However, the simulation and test require a lot of computing resources and time.
@@
Without relying on a prior knowledge, DRL learns strategies by interacting with the environment. As a result, it gains more attention in fields such as game playing[34], natural language processing[35], and robotics[12]. Recently, DRL is applied to swarm confrontation, where it learns rules from huge numbers of problem instances rather than designing them manually. De et al.[36] combine DRL with curriculum learning for pursuing an omnidirectional target with multi–agent. To improve training efficiency and learning stability, Zhang et al.[7] propose several knowledge enhancement mechanisms in a confined–space confrontation mission. However, taking the single–agent DRL approach becomes difficult faced with large–scale swarm confrontation scenarios. Therefore, Xia et al.[8] propose an end–to–end multi–agent reinforcement learning to enable agents to make decisions for cooperative target tracking. Qu et al.[37] further provide an adversarial–evolutionary game training method and designed obstacle avoidance scenarios in swarm confrontation.
@@
Among the existing methods for solving the swarm confrontation problem, expert system, game theory, and heuristic methods cannot be well applied to large–scale problems. Expert system and game theory methods are limited by prior knowledge and differential equations, respectively, and heuristic algorithms are expensive in computational resources. As a result, the performance of all three methods degrades significantly when the problem size increases. DRL is a desirable alternative due to the ability to learn strategies just by interacting with the environment. However, the direct use of an end–to–end approach in swarm confrontation [24, 25], results in the non–interpretability of the hybrid decision space, hindering the training of DRL on large–scale swarms.
@@
To deal with the above issues, the hierarchical learning method reflects the swarm confrontation to divide-–and–-conquer framework, where the commands and the actions are addressed as discrete and continue decision spaces, respectively. Hou et al.[10] integrate the finite state machine and event–condition–action frameworks to give a more interpretable solution. Wang et al.[26] decompose the swarm confrontation problem into multiple tasks to reduce the challenges of sparse reward learning. Kong et al.[38] employ the goal–conditioned HRL framework with feedback and construct a dual–aircraft formation air combat scenario. Next, Kong. [27] further extend the scenario to multiple aircraft and propose to train the lower layer through a competitive self–play approach. They make interesting attempts to introduce HRL into swarm confrontation. Inspired by them, we design a guaranteed stable HRL method for quantifying the uncertainty caused by unknown opponents’ strategies and moving obstacles. It establishes a dynamic interaction mechanism between the upper and lower layers, which has favorable potential for solving the realization problem of swarm confrontation.
@@ III-A Definition of Swarm Confrontation
This study considers the swarm confrontation as a pursuit–evasion game where exists dynamic obstacles. There are multiple heterogeneous pursuers to jointly guard the preset target point. They need to work together to capture the same number of evaders as soon as possible while avoiding collisions with obstacles and neighbors. Evaders need to get to the preset target point as quickly as possible without being caught by pursuers, as shown in Fig.1. We use blue and red agents to denote pursuers and evaders, respectively. $D_{1}$, $D_{2}$, and $D_{3}$ denote the length and width of the battlefield area and the straight–line distance between the starting areas of both sides, respectively. Let position and velocity vectors in two dimensions be denoted by $p$ and $v$, respectively. Pursuers and evaders cannot exceed the boundaries of the scenario and the maximum dynamic constraints, which can be bounded as
@@
$\lVert v\rVert\leq v_{max},$ (1)
@@
$\begin{array}[]{c}x_{ij}=\left\{\begin{array}[]{c}1,\ \text{if\ }u_{p,i}\ \text{is\ allocated\ to\ }u_{e,j}\\ 0,\ \text{otherwise}\\ \end{array}\right.,\end{array}$ (2)
where $\left[x_{ij}\right],\ i\in I,\ j\in J$ denotes the allocation matrix for all pursuers. Subsequently, pursuers chase the assigned target through real–time path planning. Throughout the process, target allocation and path planning are dynamically alternated. For pursuers and evaders, we design the rules as follows:
@@
    Both evaders and pursuers that satisfy condition 2) cannot continue moving.
@@ III-B Markov Decision Process
Let $\mathbb{R}$ denote the set of real numbers. $\mathbb{E}$ denotes the mathematical expectation. We can model reinforcement learning by a Markov decision process (MDP). We represent the MDP as a five–tuple $\left(S,A,P,R,\gamma\right)$. $S$ and $A$ denote the state of the environment and action of the agent, respectively. $P:S\times A\rightarrow S$ denotes the state transition function, $R:S\times A\rightarrow\mathbb{R}$ denotes the reward function, and $\gamma$ denotes the discount factor.
@@
IV Main Results
In this section, we introduce the guaranteed stable HRL method for solving the swarm confrontation problem in the dynamic obstacles environment. In the upper layer, it constructs an MDP model and designs a centralized deep Q–learning (DQN) algorithm for the target allocation. In the lower layer, it establishes a Dec–POMDP model and proposes a multi–agent deep deterministic policy gradient (MADDPG) algorithm for path planning. Then, the method feeds the cumulative rewards from the lower to the upper, and adopts a probabilistic ensemble network to quantify the uncertainty caused by unknown evaders’ strategies and moving obstacles. Based on the uncertainty quantification, we use an adaptive truncation method to optimize the interaction frequency between the two layers. In addition, we employ an improved model–based value estimation method to enhance the sample utilization in the upper layer which has fewer samples. Afterward, we design an integration training method including pre–training and cross–training to enhance the training efficiency and stability of our approach.
@@ IV-A Upper Layer for Target Allocation
    State. The state mainly consists of the current allocation of all pursuers $\left[x_{ij}\right]$ as well as the information of pursuer $i$ that currently needs to be allocated. It can be given as
    $s\ =\ \left(\left[x_{ij}\right],p_{i},v_{i},\rho_{c,i},v_{max,i}\right).$ (3)
@@
    Action. The action consists of the information of the selected evader, which is
    $a=\left(p_{j},v_{j},v_{max,j}\right).$ (4)
@@
    Reward. Based on the critical model of the target allocation proposed in [10], we get the one–step reward
    $r=f\left(u_{p,i},u_{e,j}\right),$ (5)
    where $f(\cdot)$ denotes the critical model with allocating $u_{e,j}$ to $u_{p,i}$.
@@
    Training method. Combining the above MDP settings, we adopt double DQN for training on target allocation. The method estimates the optimal state–action value function $Q^{*}:S\times A\rightarrow\mathbb{R}$ through a parameterized neural network $Q_{\theta}\left(s_{t},a_{t}\right)\approx Q^{*}\left(s_{t},a_{t}\right)=\mathbb{E}\left[R\left(s_{t},a_{t}\right)+\gamma\max Q^{*}\left(s_{t+1},a_{t+1}\right)\right],\forall s_{t}\in S$. The subscript $\theta$ is the weighting factor of the neural network. For $\gamma\approx 1$, $Q^{*}$ estimates the discounted returns of the optimal strategy over an infinite range. The method approximates $Q^{*}$ by $Q_{\theta}$ and the loss function is set in the following form:
    $\begin{array}[]{c}L_{\theta}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathfrak{D}}\lVert Q_{\theta}\left(s_{t},a_{t}\right)-y\rVert^{2},\end{array}$ (6)
@@
    where $\zeta\in\left[0,1\right)$ is a constant factor. $\mathfrak{D}$ is a replay buffer that iteratively grows as data are updated. We use the same DQN network with a forward inference structure as [39], thus decoupling the network from the size of the problem, which is similar to the critic network in deep deterministic policy gradient (DDPG). Based on centralized DQN, the pursuers maximize the cumulative rewards in (5) with a cooperative approach for optimal target allocation.

### v2

@@ Abstract
In swarm robotics, confrontation including the pursuit–evasion game is a key scenario. High uncertainty caused by unknown opponents’ strategies, dynamic obstacles, and insufficient training complicates the action space into a hybrid decision process. Although the deep reinforcement learning method is significant for swarm confrontation since it can handle various sizes, as an end–to–end implementation, it cannot deal with the hybrid process. Here, we propose a novel hierarchical reinforcement learning approach consisting of a target allocation layer, a path planning layer, and the underlying dynamic interaction mechanism between the two layers, which indicates the quantified uncertainty. It decouples the hybrid process into discrete allocation and continuous planning layers, with a probabilistic ensemble model to quantify the uncertainty and regulate the interaction frequency adaptively. Furthermore, to overcome the unstable training process introduced by the two layers, we design an integration training method including pre–training and cross–training, which enhances the training efficiency and stability. Experiment results in both comparison, ablation, and real–robot studies validate the effectiveness and generalization performance of our proposed approach. In our defined experiments with twenty to forty agents, the win rate of the proposed method reaches around ninety percent, outperforming other traditional methods.
@@ I Introduction
Swarm confrontation, being a typical complex problem solving, naturally consists of discrete and continuous spaces under an uncertain environment. Illustratively, commands on the battlefield always exist in the form of discrete decisions, while actions usually take place in continuous time. Recent works [24, 25] combine commands and actions into multiple high–level spaces. This method blurs the interpretability of the spaces inevitably resulting in slowly converging algorithms for large–scale swarms. Alternatively, it is not hard to reflect the swarm confrontation to divide–and–conquer framework, where the commands are translated into target allocation and the actions are addressed as path planning, from the artificial intelligence perspective[10]. Wang et al. [26] make an interesting attempt for the first time to introduce HRL into swarm confrontation. Inspired by the method, [27] decomposes multi–aircraft formation air combat into high–level strategy and sub–strategy, achieving favorable effects in confrontation among a few robots. However, the fact that we cannot guarantee the stability of the algorithms prevents us from building a bridge between HRL and swarm confrontation. Others [28, 19] have already improved the stability of the algorithm for HRL in a different scenario. They design an interactive training strategy, including pre–training, intensive training, and alternate training, to ensure the stability guaranteed. A prerequisite for this strategy is only the lower layer pre–trained, while the training of the upper is limited by the cumulative feedback from the lower. Thus we cannot directly implement it to swarm confrontation for the environmental uncertainty will make the global optimization impossible.
@@
Here, we propose a guaranteed stable HRL method, which induces quantified uncertainty into an interaction mechanism linking the allocation and planning layers, to solve the hybrid problem of swarm confrontation in various sizes and environments. Firstly, we construct two–layer DRL networks to reflect commands and actions into target allocation and path planning, respectively, since the high uncertainty caused by the nature of the confrontation, including variant opponents’ strategies and transient battlefield environment, demands a hybrid, flexible, and robust intelligent algorithm. Secondly, the mechanism, which is embedded with a probabilistic ensemble model [29] quantifying the uncertainty, regulates the interaction frequency between the two layers. The essence is the frequency is increased as the circumstance becomes uncertain and dire. Thirdly, a novel integration training method (ITM), consisting of pre–training and cross–training, ensures the stability of HRL, in which, notably, we combine pre–trained and an improved model–based value expansion (IMVE) [30] method together to fasten the convergent speed of the upper in case of few samples given by the lower. Finally, extensive experiments on different–size swarms verify that our approach outperforms the baselines including non–learning approaches and traditional DRL, and they also demonstrate the necessity of the adaptive frequency approach and ITM through ablation studies. Plus, our method shows that a trained model under a small scale holds favorable generalization in various scales of swarm confrontation. The main contributions of this paper are summarized as follows.
  • Unveiling the intrinsic mechanism of confrontation, we reflect discrete commands and continuous actions into target allocation and path planning, respectively, and then propose a novel HRL framework including discrete and continuous networks for allocation and planning.
  • We explore that high uncertainty in confrontation is an influential factor between commands and actions, so it is necessary to embed a probabilistic ensemble model by regulating the frequency adaptively for both connecting the two networks and overcoming the uncertainty.
  • Since traditional training methods may be unstable for the HRL framework in our case, we present ITM to ensure the stability of HRL, where we pre–train the commands and actions networks independently and cross–train the two networks facilitating the adaptive interaction mechanism.
@@
We organize the rest of the paper as follows. Section II introduces the research related to our study. Section III presents the formulation of our problem and provides the preliminaries on DRL. Section IV offers a detailed description and implementation of our two–layer networks and guaranteed stable HRL method. Section V describes the experiments of our method. Section VI presents our conclusions.
@@ II Related Works
The expert system method [10, 31] models the system with prior knowledge of human experts and selects the strategies in the knowledge base by fuzzy matching approach. The method relies on rules developed by an enormous number of human experts and is unable to ensure the optimality of decisions in a complex confrontation environment. In game theory[11, 32], the swarm confrontation problem is frequently modeled as a differential game. It suffers from problems such as too many state variables and complex differential equations, making it difficult to apply to complex multi–agent environments. The heuristic approach [3, 33] considers modeling the swarm as biologically inspired networks to simulate the dynamics of swarm confrontation, which has more potential to solve large–scale confrontational problems. However, the simulation and test require a lot of computing resources and time.
@@
Without relying on prior knowledge, DRL learns strategies by interacting with the environment. As a result, it gains more attention in fields such as game playing[34], natural language processing[35], and robotics[12]. Recently, DRL has been applied to swarm confrontation, where it learns rules from huge numbers of problem instances rather than designing them manually. De et al.[36] combine DRL with curriculum learning for pursuing an omnidirectional target with multi–agent. However, taking the single–agent DRL approach becomes difficult when faced with large–scale swarm confrontation scenarios. Therefore, Xia et al.[8] propose an end–to–end multi–agent reinforcement learning to enable agents to make decisions for cooperative target tracking. Qu et al.[37] further provide an adversarial–evolutionary game training method and designed obstacle avoidance scenarios in swarm confrontation.
@@
Among the existing methods for solving the swarm confrontation problem, expert system, game theory, and heuristic methods cannot be well applied to large–scale problems. The performance of all three methods degrades significantly when the problem size increases. DRL is a desirable alternative due to the ability to learn strategies just by interacting with the environment. However, the direct use of an end–to–end approach in swarm confrontation [24, 25], results in the non–interpretability of the hybrid decision space, hindering the training of DRL on large–scale swarms.
@@
To deal with the above issues, the hierarchical learning method reflects the swarm confrontation to divide-–and–-conquer framework, where the commands and the actions are addressed as discrete and continue decision spaces, respectively. Hou et al.[10] integrate the finite state machine and event–condition–action frameworks to give a more interpretable solution. Wang et al.[26] decompose the swarm confrontation problem into multiple tasks to reduce the challenges of sparse reward learning. Kong et al.[38] employ the goal–conditioned HRL framework with feedback and construct a dual–aircraft formation air combat scenario. They make interesting attempts to introduce HRL into swarm confrontation. Inspired by them, we design a guaranteed stable HRL method for quantifying the uncertainty caused by unknown opponents’ strategies, moving obstacles, and insufficient training. It establishes a dynamic interaction mechanism between the upper and lower layers, which has favorable potential for solving the realization problem of swarm confrontation.
@@ III-A Definition of Swarm Confrontation
This study considers the swarm confrontation as a pursuit–evasion game where exists dynamic obstacles. There are multiple pursuers under different abilities’ constraints to jointly guard the preset target point. They need to work together to capture evaders as soon as possible while avoiding collisions with obstacles and neighbors. Evaders need to get to the preset target point as quickly as possible without being caught by pursuers, as shown in Fig.1. We use blue and red agents to denote pursuers and evaders, respectively. $D_{1}$, $D_{2}$, and $D_{3}$ denote the length and width of the battlefield area and the straight–line distance between the starting areas of both sides, respectively. To simplify the movement of agents, we consider the pursuit–evasion game in a two–dimensional scenario. Let position and velocity vectors in two dimensions be denoted by $p$ and $v$, respectively. Pursuers and evaders cannot exceed the boundaries of the scenario and the maximum dynamic constraints, which can be bounded as
@@
$\lVert v\rVert\leq\lVert v\rVert_{\max},$ (1)
@@
Let $U_{p}=\left\{u_{p,i}|i=1,...,I\right\}$ and $U_{e}=\left\{u_{e,j}|j=1,...,J\right\}$ denote a swarm of pursuers and evaders, respectively. Both pursuers and evaders should avoid dynamic obstacles $O=\left\{o_{k}|k=1,...,K\right\}$. For any pursuer $u_{p,i}=\left(p,v,\rho_{c},\lVert v\rVert_{\max}\right)$, $\rho_{c}$ denotes the capture radius. For any evader $u_{e,j}=\left(p,v,\lVert v\rVert_{\max}\right)$, its maximum velocity magnitude is greater than the pursuers. We consider the agent and obstacle as the circle models with radius $\rho_{u}$ and $\rho_{o}$, respectively. In this paper, we design the HRL method for pursuers, while evaders use the artificial potential field method [36]. The HRL method reflects the commands and actions of the pursuit–evasion game into target allocation and path planning, respectively, through the divide–and–conquer framework approach. We can use $x_{ij}$ to denote the allocation variable which represents whether to allocate $u_{p,i}$ to $u_{e,j}$. It yields that
$\begin{array}[]{c}x_{ij}=\left\{\begin{array}[]{c}1,\ \text{if\ }u_{p,i}\ \text{is\ allocated\ to\ }u_{e,j}\\ 0,\ \text{otherwise}\\ \end{array}\right..\end{array}$ (2)
$\left[x_{ij}\right],\ i\in I,\ j\in J$ denotes the allocation matrix for all pursuers. Subsequently, pursuers chase the assigned target through real–time path planning. Throughout the process, target allocation and path planning are dynamically alternated. For pursuers and evaders, we design the rules as follows:
@@
    Both evaders and pursuers that satisfy rule 2) cannot continue moving.
@@
Rule 3) explains that a pursuer can capture up to one evader. We set up that there is the same number of pursuers and evaders in the pursuit–evasion game. If pursuers are less than evaders, pursuers cannot capture all the evaders in this setting. Otherwise, if pursuers are more than evaders, it will be not critical for pursuers to develop a strategy to capture as many evaders as possible.
@@ III-B Markov Decision Process
Let $\mathbb{R}$ denote the set of real numbers. $\mathbb{E}$ denotes the mathematical expectation. We can model reinforcement learning by a Markov decision process (MDP). We represent the MDP as a five–tuple $\left(S,A,P,R,\gamma\right)$. $S$ and $A$ denote the state of the environment and action of the agent, respectively. $P\left(s,a\right):S\times A\times S\rightarrow[0,1]$ denotes the state transition probabilities, $R:S\times A\rightarrow\mathbb{R}$ denotes the reward function, and $\gamma$ denotes the discount factor.
@@
IV Methodology
In this section, we introduce the guaranteed stable HRL method for solving the swarm confrontation problem in the dynamic obstacles environment. In the upper layer, it constructs an MDP model and designs a centralized deep Q–learning (DQN) algorithm for the target allocation. In the lower layer, it establishes a Dec–POMDP model and proposes a multi–agent deep deterministic policy gradient (MADDPG) algorithm for path planning. Then, the method feeds the cumulative rewards from the lower to the upper, and adopts a probabilistic ensemble network to quantify the uncertainty caused by unknown evaders’ strategies, moving obstacles, and insufficient training. Based on the uncertainty quantification, we use an adaptive truncation method to optimize the interaction frequency between the two layers. In addition, we employ an improved model–based value expansion method to enhance the sample utilization in the upper layer which has fewer samples. Afterward, we design an integration training method including pre–training and cross–training to enhance the training efficiency and stability of our approach.
@@ IV-A Upper Layer for Target Allocation
    State. The state mainly consists of the current allocation of all pursuers $\left[x_{ij}\right]$ as well as information on pursuer $i$ that currently needs to be allocated. It can be given as
    $s=\left(\left[x_{ij}\right],p_{i},v_{i},\rho_{c,i},\lVert v\rVert_{\max,i}\right).$ (3)
@@
    Action. The action consists of information on the selected evader, which is
    $a=\left(p_{j},v_{j},\lVert v\rVert_{\max,j}\right).$ (4)
@@
    Reward. The capture probability of pursuer $i$ relative to evader $j$ can be assessed by the distance between them and the capture radius, which is described as
    $q_{ij}=\frac{\rho_{c,i}}{\rho_{c,i}+\lVert p_{i}-p_{j}\rVert}.$ (5)
    If there are multiple pursuers assigned to the same evader $j$, then the joint capture probability is
    $\overline{q}_{j}=1-\prod_{i}{\left(1-q_{ij}\right),~{}i}\in\left\{k\left|x_{kj}=1\right.\right\}.$ (6)
    The capture value of the evader is related to its maximum velocity magnitude, therefore, the goal of target allocation can be set to maximize the following effectiveness function:
    $\mathbb{E}\left(\left[x_{ij}\right]\right)$ $=\sum_{1}^{J}{\overline{q}_{j}\lVert v\rVert_{\max,j}}$
    $=\sum_{1}^{J}{\left[1-\prod_{i}{\left(1-q_{ij}\right)}\right]\lVert v\rVert_{\max,j}}.$ (7)
    Let $\left[\tilde{x}_{ij}\right]$ denote the allocation matrix after allocating pursuer $i$ to evader $j$. The local reward in allocation can be obtained as
    $r_{\text{allo},1}=\mathbb{E}\left(\left[\tilde{x}_{ij}\right]\right)-\mathbb{E}\left(\left[x_{ij}\right]\right).$ (8)
    Once all the pursuers have completed their assignments, we get the global reward in allocation, which can be calculated as
    $r_{\text{allo},2}=\frac{1}{I}\mathbb{E}\left(\left[x_{f}\right]\right),$ (9)
    where $\left[x_{f}\right]$ denotes the allocation matrix in final. The static reward for target allocation is a linear operation consisting of the local and global allocation rewards, which can be expressed as follows:
    $r_{\text{allo}}=\omega_{1}r_{\text{allo},1}+\left(1-\omega_{1}\right)r_{\text{allo},2},$ (10)
    where $\omega_{1}$ is the weighting factor between the local and global allocation rewards.
@@
    Training method. Combining the above MDP settings, we adopt double DQN for training on target allocation. The method estimates the optimal state–action value function $Q^{*}:S\times A\rightarrow\mathbb{R}$ through a parameterized neural network $Q_{\theta}\left(s_{t},a_{t}\right)\approx Q^{*}\left(s_{t},a_{t}\right)=\mathbb{E}\left[R\left(s_{t},a_{t}\right)+\gamma\max Q^{*}\left(s_{t+1},a_{t+1}\right)\right],\forall s_{t}\in S$. The subscript $\theta$ is the weighting factor of the value–function network. For $\gamma\approx 1$, $Q^{*}$ estimates the discounted returns of the optimal strategy over an infinite range. The method approximates $Q^{*}$ by $Q_{\theta}$ and the loss function is set in the following form:
    $\begin{array}[]{c}L_{\theta}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathcal{D}}\lVert Q_{\theta}\left(s_{t},a_{t}\right)-y\rVert^{2},\end{array}$ (11)
@@
    where $\zeta\in\left[0,1\right)$ is a constant factor. $\mathcal{D}$ is a replay buffer that iteratively grows as data are updated. We use the same DQN network with a forward inference structure as [39], thus decoupling the network from the size of the problem, which is similar to the critic network in deep deterministic policy gradient (DDPG). Based on centralized DQN, the pursuers maximize the cumulative rewards in (10) with a cooperative approach for optimal target allocation.